OpenStudy (anonymous):

Using the Pythagorean Triple what is the missing side length if A=5.6 an B=4.5.

OpenStudy (apoorvk):

let missing length= k then k^2 = a^2 + b^2 (according to pythagoras theorem) substitute and evaluate!!!

OpenStudy (anonymous):

5.6^2 + 4.5^2 = 51.61 then take root51.61 to get 7.184